OpenStudy (anonymous):

A rectangular table top has a length of 5xcm and a width of xcm. An expression for it's perimter is?

OpenStudy (lgbasallote):

P = 2(length + width) so P = 2(5x + x) P = 2 (6x) P = 12x

OpenStudy (unklerhaukus):